# Deep learning approach for identification of Hii regions during reionization in 21-cm observations – III. image recovery

### v1

@@ Acknowledgements
The authors would like to thank Andrei Mesinger and Adrian Liu for their useful discussions and comments. The authors also would like the organizers of the Cosmic Dawn at High Latitude Nordita workshop. MB acknowledges the financial support from the Swiss National Science Foundation (SNSF) under the Sinergia Astrosignals grant (CRSII5_193826). We acknowledge access to Piz Daint at the Swiss National Supercomputing Centre, Switzerland, under the SKA’s share with the project ID sk014. This work has been done in partnership with the SKACH consortium through SERI funding. Nordita is supported in part by NordForsk.
@@
The data underlying this article is available upon request and can also be re-generated from scratch using the publicly available 21cmFAST (Mesinger et al., 2011; Murray et al., 2020) and Tools21cm (Giri et al., 2020) code. The SERENEt code and its trained network weights are available on the author’s GitHub page: https://github.com/micbia/serenet.
@@ Appendix A Application to Large Images
Though SERENEt has been trained on 2D image data with a mesh size of $128$ per side, we can easily apply it to images with larger mesh dimensions. In Figure 14, we show an example applied on an image with $1024^{2}$ pixels. We apply a kernel with an overlapping factor of $\frac{1}{32}$. This factor indicates the portion of the image that overlaps. For instance, a factor of $1$ is the case where each kernel is applied back-to-back with no overlap, while $\frac{1}{2}$ is the case where half the image is overlapped.
@@
The inset plot shows a zoom-in on the bottom left corner of the larger image. The zoom-in size corresponds to the kernel dimensions and the shape of the SERENEt input, as mentioned in §3.1. As we can see from the image, the pixels on the borders of the image are under-sampled, and only the very first $4\times 4$ pixels are sampled only once. However, from the SKA-Low tomographic image with $\rm{FoV}\simeq 10\,\rm deg$, we expect that only the central $4\times 4\,\rm deg$ region will be processed for summary statistic analysis.
@@
Because of the selected overlapping factor, this setup virtually creates 50 1765017650\,17650 176 of partially overlapping images that SERENEt need to process and eventually puzzle back together. Overlapping pixels are combined with a simple average based on the number of overlaps, indicated by the colorbar in Figure 14.
@@
Thanks to the sliding overlapping kernel approach, SERENEt can be applied to images with much larger mesh sizes and avoid requiring to re-train the network, saving computational time.
@@
Figure 14: Example of SERENEt applied to an image with a large FoV. The inset plot shows a region with a mesh size of $128^{2}$ and comoving size of $1.56\,\rm Mpc$ located at the bottom left corner. This dimension corresponds to the input shape of our networks. The colorbar indicates the number of times the overlapping kernel samples each pixel.

### v2

@@ Acknowledgements
The authors would like to thank Andrei Mesinger and Adrian Liu for their valuable discussions and comments. The authors would also like to thank the organizers of the "Cosmic Dawn at High Latitude" Nordita workshop program for providing a conducive platform to develop this project further. MB acknowledges the financial support from the Swiss National Science Foundation (SNSF) under the Sinergia Astrosignals grant (CRSII5_193826). We acknowledge access to Piz Daint at the Swiss National Supercomputing Centre, Switzerland, under the SKA’s share with the project ID sk014. This work has been done in partnership with the SKACH consortium through SERI funding. Nordita is supported in part by NordForsk.
@@
The data underlying this article is available upon request and can also be regenerated from scratch using the publicly available 21cmFAST (Mesinger et al., 2011; Murray et al., 2020) and Tools21cm (Giri et al., 2020) code. The SERENEt code and its trained network weights are available on the author’s GitHub page: https://github.com/micbia/serenet.
@@ Appendix A Application to Large Images
Although SERENEt has been trained on 2D image data with a mesh size of $128$ per side, it can be easily applied to images with larger mesh dimensions. In Figure 14, we show an example applied to an image with $1024^{2}$ pixels. We use a kernel with an overlapping factor of $\frac{1}{32}$. This factor indicates the portion of the image that overlaps. For instance, a factor of $1$ is the case where each kernel is applied back-to-back with no overlap, while $\frac{1}{2}$ is the case where half the image is overlapped.
@@
The inset plot shows a zoom-in on the bottom left corner of the larger image. The zoom-in size corresponds to the kernel dimensions and the shape of the SERENEt input, as mentioned in §3.1. As we can see from the image, the pixels on the borders of the image are undersampled, and only the very first $4\times 4$ pixels are sampled only once. However, from the SKA-Low tomographic image with $\rm{FoV}\simeq 10\,\rm deg$, we expect that only the central $4\times 4\,\rm deg$ region will be processed for summary statistic analysis.
@@
Because of the selected overlapping factor, this setup virtually creates 50 17650\,176 of partially overlapping images that SERENEt need to process and eventually puzzle back together. Overlapping pixels are combined with a simple average based on the number of overlaps, indicated by the colorbar in Figure 14.
@@
Thanks to the sliding overlapping kernel approach, SERENEt can be applied to images with much larger mesh sizes, thereby avoiding the need for retraining the network and saving computational time.
@@
Figure 14: Example of SERENEt applied to an image with a large FoV. The inset plot shows a region with a mesh size of $128^{2}$ and a comoving size of $1.56\,\rm Mpc$ located at the bottom left corner. This dimension corresponds to the input shape of our networks. The colorbar indicates the number of times the overlapping kernel samples each pixel.
Appendix B The Size of the Dataset
Figure 15: Comparison of the $R^{2}$ score on the validation dataset. The dark blue line indicates the model employed in the results of this paper, which was trained on 10,000 samples. A grey line indicates the learning rate. The blue shadow indicates the corresponding range for models trained on a dataset with one-tenth of the original samples.
[…]
Here, we want to analyse and evaluate the required dataset size for a stable and accurate recovery of the 21-cm signal with U-Nets in foreground-contaminated images. Therefore, we trained a few additional models on a fraction of the entire dataset for the realistic prior scenario. For each independent training session, we randomly select 1,000 samples, corresponding to one-tenth of the whole dataset, and train the model until convergence and early stopping is reached.
From Figure 15, we show the $R^{2}$ score, Equation 3, on the validation set for these two scenarios. The dark blue line indicates the score for the model trained on the entire dataset, which consists of 10’000 samples, presented in §2.1 and used for the results of this paper, §4. In this case, we can notice that the score quickly rises after a few epochs and stabilises at $R^{2}\simeq 80\%$, reaching its best performance after approximately $55$ epochs. On the other hand, the blue shadow region indicates the corresponding score for the models trained on a reduced dataset. We can notice that for this scenario, the score range is lower, with almost $10\%$ less accuracy. Moreover, the best performance is reached only after $\approx 120$ epochs with an averaged best score around $R^{2}\simeq 70\%$.
The reason for this trend could be related to the fact that the limits of the parameter space we employ, $(\zeta,M_{\mathrm{min}},R_{\mathrm{mfp}})$, are wide, providing a vast range of possible combinations. A similar conclusion was reached in a completely different domain but with similar challenges (e.g.: Althnian et al., 2021; Dawson et al., 2023). Therefore, we conclude that training our network on only a few thousand realisations of the residual foreground contaminated lightcone is insufficient to sample the astrophysical parameter space we employ and generalise the recovery of the 21-cm signal. Moreover, our evaluation suggests that decreasing the range of the astrophysical/cosmological parameter space is necessary to avoid underperformance due to smaller datasets.